# of Diffusion Models

### v3

@@ Introduction
Table 1 summarizes our main quantitative findings. DogFit consistently performs comparatively or superior to prior methods in both FID and $\text{FD}_{\text{DINOV2}}$, while maintaining the efficiency of one-pass sampling. Unlike CFG and MG, DogFit remains effective in unconditional settings like FFHQ due to its class-agnostic domain-alignment direction. On SiT, CFG and MG slightly outperform DogFit in FID, but DogFit outperforms both CFG and MG on $\text{FD}_{\text{DINOV2}}$, underscoring the value of utilizing diverse metrics for comprehensive evaluation. Importantly, equipping DogFit with controllable guidance introduces no performance degradation while increasing the parameter count by only 1.33 million parameters (less than 2% of the initial network parameters), leaving no notable impact on sampling TFLOPS. Additional results on Precision and Recall are provided in Appendix B. Figure 4 presents a visual comparison of generated samples across target domains on DiT. Guidance-based methods consistently outperform naïve fine-tuning by producing more realistic and coherent images. DogFit generates sharp, domain-relevant details on par with or better looking than DoG, while avoiding off-distribution artifacts occasionally seen with CFG and MG (e.g., distorted faces or tires). SiT qualitative results are provided in Appendix B.
@@
Controllable Guidance
We study the effect of $\lambda$ for sampling the guidance scalar $w$ from the SED during training (Equation 8). As shown in Figure 5 (left), sparse exposure to higher guidance is sufficient for generalization and achieving a diversity-fidelity tradeoff. In contrast, excessive exposure to large guidance can harm the model’s ability to retain pretrained features, leading to instability and collapsed generations. Based on these findings, we adopt $\lambda=3$ (95% of sampled $w$s in $[1,2]$) for all experiments. Figure 5 (right) illustrates how varying guidance strength at test time effectively modulates the trade-off between fidelity and diversity, increasing sample quality while reducing variation with higher guidance values. For more analysis, see Appendix B.
Guidance Scheduling.
We analyze the impact of late-start and cut-off thresholds on generation quality and diversity. Figure 6(a) shows how varying the late-start threshold $\tau_{s}$ impacts FID, Precision, and Recall. Delaying guidance improves FID and especially Precision at first, as early stabilization allows formation of more accurate directional cues. However, setting $\tau_{s}$ too high leaves insufficient time for the model to internalize guidance, reducing it to naïve fine-tuning and lowering precision. We set $\tau_{s}=12\text{k}$ for FID and $\tau_{s}=6\text{k}$ for $\text{FD}_{\text{DINOV2}}$ in all experiments. Figure 6(b) shows the effect of cut-off threshold $\tau_{c}$ on FID, Precision, and Recall. Delaying guidance to later denoising steps enhances diversity, particularly for target datasets with strong global structural similarity to the source domain, such as FFHQ, Cars, and Caltech. However, setting $\tau_{c}$ too low limits exposure to guidance, again reducing the model to naïve fine-tuning. While the optimal $\tau_{c}$ may depend on the source–target similarity, we use $\tau_{c}=0.5$ for FID and $\tau_{c}=1$ for $\text{FD}_{\text{DINOV2}}$ in our experiments ($t\in[0,1]$). See analysis on $\text{FD}_{\text{DINOV2}}$ in Appendix B.
@@
Domain-guided Parameter-efficient Fine-tuning.
We demonstrate that DogFit is fully compatible with DiffFit, a leading PEFT method for DiT models (Xie et al. 2023). As shown in Table 3, combining the two yields high-quality generations with lower cost in both training and test-time. This highlights DogFit’s practicality as a simple drop-in enhancement for efficient fine-tuning mechanisms.
Sampling Steps.
Table 2 reports the effect of varying the number of sampling steps during generation. DogFit outperforms MG in nearly all settings, with especially notable gains at lower step counts. This underscores the effectiveness of DogFit’s target alignment in low-resource scenarios where fast sampling is critical.
@@ Conclusion
Limitations and Future Work.
While effective, the proposed guidance scheduling strategies rely on manually chosen, fixed thresholds. Future work could explore adaptive or data-driven approaches that respond to training dynamics or domain-specific properties. Moreover, although DogFit performs well on class-conditioned image domains, extending it to broader settings, i.e., text-to-image models, video generation, and other modalities, remains an exciting direction.
@@ Appendix A Additional Methodology Explanations
Figure 7 visualizes the Cumulative Distribution Function (CDF) of $w$s sampled during training from the shifted exponentially decaying distribution (SEDD) in Equation 8. Larger values of $\lambda$ concentrate the sampling mass closer to $w=1$, limiting exposure to higher guidance values during training. This design ensures that the model primarily learns under weak to moderate guidance, while still sparsely encountering stronger guidance values.
@@
General Assumptions. We assume that both the pretrained source model $\epsilon_{\theta_{0}}$ and the fine-tuned model $\epsilon_{\theta}$ share the same architecture and noise schedule $(\alpha_{t})$, and are trained to approximate the same family of score functions under a common sampling procedure.
@@
Traditionally, CFG estimates both the class-conditional (Equation 12) and the marginal (Equation 13) noises with the same diffusion model. With enough data in the target domain, the class-conditional noise can be well estimated during fine-tuning via $\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})$. However, learning the target noise estimator $\epsilon_{\theta}(x_{t}\mid\mathcal{D}^{\mathcal{T}})$ faces the risk of undertraining due to only being trained on a small portion of the small target dataset. Since the source model is usually trained on large, diverse datasets, it likely captures a good approximation of the general image manifold. Furthermore, it has been shown that diffusion models have generalizability on other data distributions (Deja et al. 2022) and can denoise noisy images from domains other than those they have been trained on (Meng et al. 2021). Therefore, we can assume that the pre-trained source model captures a general enough prior, giving us the ability to approximate the unconditional score using the pre-trained source model, leading to the formulation of DoG:
@@
which exactly matches the offset used in DoG sampling (Zhong et al. 2025). Our method takes a step further from this and applies the same guidance offset online. We train a diffusion model to directly predict this guided score instead of separately learning Equations 12 and 13 in the form of DoG.
@@
Note on the assumptions: Prior work on CFG has shown that the linearity assumption holds in practice for low to moderate $w$ values (Zheng and Lan 2023). Furthermore, our use of an SED to sample $w$ during training (see Section Incorporating Controllability) helps maintain accurate estimation of $\epsilon$ in the unconditional setting. Based on this, we conjecture that both assumptions can approximately hold in practice.
@@
Since the training loss is a squared error between two linear functions of $w$, minimizing this loss for all values of $w$ is equivalent to matching the corresponding coefficients of both functions. That is, to minimize the loss over the full support of $w$, the model must satisfy:
@@
$\epsilon_{\theta}(x_{t}\mid c,w,\mathcal{D}^{\mathcal{T}})=\epsilon+(w-1)\cdot\left(\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})-\epsilon_{\theta_{0}}(x_{t})\right),$ (21)
which is exactly the DoG-guided prediction used at test time. Therefore, DogFit learns to internalize the DoG guidance signal during training, removing the need for external guidance at sampling. $\blacksquare$
@@
This enables DoGFit to retain MG’s class-conditional guidance while avoiding out-of-distribution generation by pulling samples toward the core of the target domain manifold.
@@
According to Equations 12 and 13, Equation 24 can be rewritten as:
$\sigma(w-1)\cdot\text{sg}\left(\nabla_{x_{t}}\log p_{\theta}(x_{t}|\mathcal{D^{T}})-\nabla_{x_{t}}\log p_{\theta}(x_{t})\right)$ (25)
@@
Thus, given that MG uses $\epsilon_{\theta}(x_{t}|\mathcal{D^{T}})$ as an approximation to $\sigma_{t}\nabla_{x_{t}}\log p(x_{t}|\mathcal{D^{T}})$, we can express DogFit as applying MG, but with a correction term that pushes towards a target domain classifier’s gradient direction. $\blacksquare$
@@ Appendix B Additional Experiments
We further examine the sensitivity of our method to $w$ in the absence of controllability. We train a specific model for each $w$ and compare the FID trends of DogFit and DoG across varying values (Figure 10). While both methods achieve optimal FID in a similar range, DogFit exhibits a sharper degradation for larger $w$, suggesting higher sensitivity to strong guidance values. We hypothesize that this is due to training-time overexposure to extreme guidance, which can lead the model to learn from unnatural signals that do not follow the real target distribution, thereby reducing generation quality. This motivates our design choice in Section Incorporating Controllability to sample $w$ from an exponentially decaying distribution, ensuring the model is only sparsely exposed to high guidance values during training.
@@
Figure 8 provides an extended version of the controllable guidance analysis for the Stanford-Cars dataset. This complements the main findings and illustrates the same trends on a new domain.
Guidance Scheduling on $\text{FD}_{\text{DINOV2}}$
We observe that delaying guidance using the late-start strategy improves both $\text{FD}_{\text{DINOV2}}$ and precision, consistent with our FID-based findings. This supports the idea that postponing guidance allows for stronger and more accurate directional signals once the model has sufficiently stabilized (Figure 9a). However, unlike FID, applying the cut-off strategy—where guidance is limited to later denoising steps—shows little to no improvement in FDDINOV2 across most domains (Figure 9 (b)). We believe this is due to a fundamental difference in what the two metrics capture: FID benefits from preserving the global structure inherited from the source model, which is often formed early in the denoising process. In contrast, FDDINOV2 focuses more on semantic alignment with the target domain, which may require guidance to be present throughout the generation, not just in the later steps. This distinction further highlights the importance of using multiple metrics when evaluating generative models.
Qualitative Comparison on SiT-XL/2
Figure 11 provides a qualitative comparison on the SiT-XL/2 backbone. DogFit generates sharp, domain-relevant details on par with DoG, while avoiding out-of-distribution behavior occasionally seen with CFG and MG (e.g., distorted faces and birds).
Test-time Uniform SEDD
$\mathbf{\omega}$ $\mathcal{U}[1,1.1]$ $\mathcal{U}[1,1.25]$ $\mathcal{U}[1,1.5]$ $\mathcal{U}[1,2]$ ($\mathbf{\lambda=2}$)
@@
Precision and Recall Analysis
Table 9 presents a comparison of Precision and Recall (Kynkäänniemi et al. 2019) between DogFit and the baselines on the DiT-XL/2 backbone. Precision reflects the fidelity or quality of generated samples, while Recall measures sample diversity. As expected, all guidance methods consistently improve Precision compared to finetuning while reducing Recall, aligning with the general behavior of guided diffusion models. This effect arises because guidance steers sampling toward high-likelihood regions of the conditional distribution, which inherently narrows the output distribution and suppresses diversity. Our method, both with and without guidance control, matches the performance of DoG in terms of both Precision and Recall in most settings. Notably, DogFit achieves the highest or second-highest Precision in both Inception and DINOV2 metrics across all datasets). In terms of Recall, DogFit and DoG slightly underperform CFG and MG, which we believe is due to the additional domain alignment term further reducing the diversity to prevent out-of-domain target generations (Equation 22). While recent works explore diversity-preserving techniques and scheduling mechanisms for guidance during sampling (Kynkäänniemi et al. 2024; Sadat et al. 2023), analogous techniques for training-time guidance have yet to be developed. We leave the exploration of such methods for training-guided models like DogFit as an exciting direction for future work.
@@
Table 5 provides a comparison between DogFit and CFG when fine-tuning DiT-XL/2 pretrained at $512{\times}512$ on Food101 (Bossard, Guillaumin, and Van Gool 2014) at $512{\times}512$. Results indicate that dogFit is fully scalable to larger resolutions, broadening its applicability and demonstrating invariance to image scale. For this experiment, we use a batch size of 12 and use the same hyperparameters as in our main experiments.
Robustness under Stronger Domain Shifts
While our target datasets span a range of distribution gaps, we additionally evaluate transfer on the EuroSat dataset (helber2019eurosat), consisting of 27,000 satelite images with 10 classes, which shares very few features with the source domain. This setting probes whether DogFit’s source prior still provides a better signal than that of the fine-tuned target under extreme shift. We note that as the target domain diverges substantially from the source, the advantages of transfer learning from the source generally diminishes relative to training on the target data from scratch. Nevertheless, our results in Table 6 show that DogFit still outperforms Fine-tuning and CFG in this case, indicating that leveraging the source prior can still improve generation quality even when transfer benefits are attenuated.
@@
Table 6 shows that sampling the guidance weight $\omega$ from a SEDD (see Figure 7) consistently outperforms uniform schedules across all test-time values of $\omega$. In particular, SEDD yields lower FID while uniform sampling over broader ranges destabilizes learning and collapses diversity, with FID inflating as $\omega$ increases. This empirical trend aligns with our linearity assumption in Proposition 1 which leads to the equivalence of DogFit and DoG (Zhong et al. 2025) under small $\omega$ values (Zheng and Lan 2023). By concentrating the probability mass near $\omega\!\approx\!1$, SEDD keeps optimization within the proposition’s validity region, preserving the expected guidance behavior and allowing the use of larger test-time guidance values during generation.
Applying DogFit to Text-to-Image transfer learning
To further show the broad applicability of our method, we show the efficacy of DogFit in two different text-to-image experiments, Dreambooth subject driven transfer  (ruiz2023dreambooth), and LoRA style transfer  (hu2022lora) to Pokémon. Note that these settings typically require large guidance values of up to $7.5$ for the model to produce visually pleasing images. However, DogFit is more theoretically grounded when exposed to small-guidance values during training, where it locally matches DoG (Zhong et al. 2025) (see Proposition 1). While controllable guidance (conditioning on $w$) is a natural extension allowing for large test-time $\omega$ values, we focus here on fixed guidance during training due to limited time and leave that extension for future work. Despite this conservative setting, DogFit significantly improves unguided fine-tuning with small guidance values ($\omega=[1.65,1.75]$) and approaches the performance of CFG with high guidance value ($\omega=7.5$).
@@
We apply Dreambooth fine-tuning on Stable Diffusion v1.5 (Rombach et al. 2022; podell2023sdxl) to new subjects using a handful of instance images with DogFit. In this regime, the fine-tuning prompts are typically limited to templates such as “a [v] photo of {subject}, while at inference we expect the model to respond to diverse prompts from the source domain. This mismatch makes faithful distillation of text-conditioning challenging. Nevertheless, Fig. 12 and Tab. 7 show that DogFit learns controllable alignment with the prompt and significantly outperforms unguided fine-tuning while approaching CFG on DINO, CLIP-I, and CLIP-T, the standard metrics for subject-driven generation (ruiz2023dreambooth).
@@
We adapt Stable Diffusion v1.5 to the Pokémon style dataset [^2] https://huggingface.co/datasets/diffusers/pokemon-gpt4-captions with 833 image–caption pairs. We train a LoRA adapter (rank 4) for 2,000 iterations at a learning rate of $10^{-4}$. We evaluate with CLIP-T over the first 100 prompts (one image per prompt). As summarized in Table 8, DogFit significantly improves prompt alignment over unguided fine-tuning while using a small guidance value ($\omega{=}1.75$) and a single sampling pass, and it closely approaches CFG which requires two passes and a large guidance value ($\omega{=}7.5$).
Metric Method Unlabelled Labelled
FFHQ ArtBench Caltech CUB-Birds Food Stanford-Cars Avg.
Precision (Inception) Fine-tuning 0.70 0.67 0.60 0.58 0.79 0.51 0.63
+ CFG (Ho and Salimans 2022) - 0.68 0.76 0.72 0.82 0.63 0.72
+ DoG (Zhong et al. 2025) 0.83 0.77 0.81 0.82 0.88 0.77 0.81
MG (Tang et al. 2025) - 0.68 0.76 0.71 0.81 0.62 0.72
DogFit 0.81 0.76 0.75 0.82 0.87 0.74 0.79
DogFit + Control 0.83 0.76 0.78 0.83 0.87 0.80 0.81
Precision (DINOV2) Fine-tuning 0.74 0.84 0.48 0.16 0.41 0.30 0.44
+ CFG (Ho and Salimans 2022) - 0.83 0.67 0.30 0.48 0.47 0.52
+ DoG (Zhong et al. 2025) 0.82 0.91 0.73 0.39 0.59 0.61 0.65
MG (Tang et al. 2025) - 0.84 0.63 0.28 0.51 0.44 0.54
DogFit 0.83 0.92 0.67 0.38 0.58 0.58 0.62
DogFit + Control 0.86 0.90 0.67 0.38 0.58 0.59 0.62
Recall (Inception) Fine-tuning 0.67 0.57 0.75 0.81 0.52 0.58 0.65
+ CFG (Ho and Salimans 2022) - 0.56 0.66 0.76 0.53 0.59 0.62
+ DoG (Zhong et al. 2025) 0.61 0.50 0.65 0.68 0.48 0.52 0.57
MG (Tang et al. 2025) - 0.56 0.67 0.75 0.55 0.59 0.624
DogFit 0.63 0.51 0.70 0.68 0.50 0.55 0.59
DogFit + Control 0.57 0.51 0.68 0.67 0.50 0.48 0.57
Recall (DINOV2) Fine-tuning 0.46 0.21 0.66 0.95 0.67 0.82 0.66
+ CFG (Ho and Salimans 2022) - 0.22 0.59 0.87 0.61 0.84 0.63
+ DoG (Zhong et al. 2025) 0.44 0.21 0.57 0.83 0.64 0.83 0.62
MG (Tang et al. 2025) - 0.22 0.60 0.89 0.60 0.84 0.63
DogFit 0.43 0.21 0.60 0.83 0.64 0.83 0.62
DogFit + Control 0.43 0.21 0.61 0.85 0.64 0.83 0.63

### v4

@@ Introduction
Tab.1 summarizes our main quantitative findings. DogFit consistently performs comparatively or superior to prior methods in both FID and $\text{FD}_{\text{DINOV2}}$, while maintaining the efficiency of one-pass sampling. Unlike CFG and MG, DogFit remains effective in unconditional settings like FFHQ due to its class-agnostic domain-alignment direction. On SiT, CFG and MG slightly outperform DogFit in FID, but DogFit outperforms both CFG and MG on $\text{FD}_{\text{DINOV2}}$, underscoring the value of utilizing diverse metrics for comprehensive evaluation. Importantly, equipping DogFit with controllable guidance introduces no performance degradation while increasing the parameter count by only 1.33 million parameters (less than 2% of the initial network parameters), leaving no notable impact on sampling TFLOPS. Additional results on Precision and Recall are provided in Appx. B. Fig. 4 presents a visual comparison of generated samples across target domains on DiT. Guidance-based methods consistently outperform naïve fine-tuning by producing more realistic and coherent images. DogFit generates sharp, domain-relevant details on par with or better looking than DoG, while avoiding off-distribution artifacts occasionally seen with CFG and MG (e.g., distorted faces or tires). SiT qualitative results are provided in Appx. B.
@@
Controllable Guidance We study the effect of $\lambda$ for sampling the guidance scalar $w$ from the SED during training (Eq. 8). As shown in Fig. 5 (left), sparse exposure to higher guidance is sufficient for generalization and achieving a diversity-fidelity tradeoff. In contrast, excessive exposure to large guidance can harm the model’s ability to retain pretrained features, leading to instability and collapsed generations. Based on these findings, we adopt $\lambda=3$ (95% of sampled $w$s in $[1,2]$) for all experiments. Fig. 5 (right) illustrates how varying guidance strength at test time effectively modulates the trade-off between fidelity and diversity, increasing sample quality while reducing variation with higher guidance values (more analysis in Appx. B).
Guidance Scheduling. We analyze the impact of late-start and cut-off on generation quality and diversity. Fig. 6(a) shows how varying the late-start threshold $\tau_{s}$ impacts FID, Precision, and Recall. Delaying guidance improves FID and especially Precision at first, as early stabilization allows formation of more accurate directional cues. However, setting $\tau_{s}$ too high leaves insufficient time for the model to internalize guidance, reducing it to naïve fine-tuning and lowering precision. We set $\tau_{s}=12\text{k}$ for FID and $\tau_{s}=6\text{k}$ for $\text{FD}_{\text{DINOV2}}$ in all experiments. Fig. 6(b) shows the effect of cut-off threshold $\tau_{c}$ on FID, Precision, and Recall. Delaying guidance to later denoising steps enhances diversity, particularly for target datasets with strong global structural similarity to the source domain, such as FFHQ, Cars, and Caltech. However, setting $\tau_{c}$ too low limits exposure to guidance, again reducing the model to naïve fine-tuning. While the optimal $\tau_{c}$ may depend on the source–target similarity, we use $\tau_{c}=0.5$ for FID and $\tau_{c}=1$ for $\text{FD}_{\text{DINOV2}}$ in our experiments ($t\in[0,1]$). See $\text{FD}_{\text{DINOV2}}$ analysis in Appx. B.
@@
Domain-guided Parameter-efficient Fine-tuning. We demonstrate that DogFit is fully compatible with DiffFit, a leading PEFT method for DiT models (xie2023difffit). As shown in Tab.3, combining the two yields high-quality generations with lower cost in both training and test-time. This highlights DogFit’s practicality as a simple drop-in enhancement for efficient fine-tuning mechanisms.
Sampling Steps. Tab.2 reports the effect of varying the number of sampling steps during generation. DogFit outperforms MG in nearly all settings, with especially notable gains at lower step counts. This underscores the effectiveness of DogFit’s target alignment in low-resource scenarios where fast sampling is critical.
@@ Conclusion
Limitations and Future Work. While effective, the proposed guidance scheduling strategies rely on manually chosen, fixed thresholds. Future work explores adaptive or data-driven approaches that respond to training dynamics or domain-specific properties. Moreover, extension of DogFit it to broader settings, i.e., text-to-image, audio, and video generation remains an exciting direction.
Acknowledgments. This research was supported by the Natural Sciences and Engineering Research Council of Canada, and the Digital Research Alliance of Canada.
@@ Appendix A Additional Methodology Explanations
Fig. 7 visualizes the Cumulative Distribution Function (CDF) of $w$s sampled during training from the shifted exponentially decaying distribution (SEDD) in Equation 8. Larger values of $\lambda$ concentrate the sampling mass closer to $w=1$, limiting exposure to higher guidance values during training. This design ensures that the model primarily learns under weak to moderate guidance, while still sparsely encountering stronger guidance values.
@@
General Assumptions. We assume that both the pretrained source model $\epsilon_{\theta_{0}}$ and the fine-tuned model $\epsilon_{\theta}$ share the same architecture and noise schedule $\alpha_{t}$, and are trained to approximate the same family of score functions under a common sampling procedure.
@@
Traditionally, CFG estimates both the class-conditional (Eq. 12) and the marginal (Eq. 13) noises with the same diffusion model. With enough data in the target domain, the class-conditional noise can be well estimated during fine-tuning via $\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})$. However, learning the target noise estimator $\epsilon_{\theta}(x_{t}\mid\mathcal{D}^{\mathcal{T}})$ faces the risk of under-training due to only being trained on a small portion of the tiny target dataset (typically 10%  (ho2022classifier). Since the source model is usually trained on large, diverse datasets, it likely captures a good approximation of the general image manifold. Furthermore, it has been shown that diffusion models have generalizability on other data distributions (deja2022analyzing) and can denoise noisy images from domains other than those they have been trained on (meng2021sdedit). Therefore, we can assume that the pre-trained source model captures a general enough prior, giving us the ability to approximate the unconditional score using the pre-trained source model. This leads us to the formulation of DoG:
@@
which exactly matches the offset used in DoG sampling (zhong2025domain). Our method takes a step further from this and applies the same guidance offset online. We train a diffusion model to directly predict this guided score instead of separately learning Eq. 12 and Eq. 13 in the form of DoG.
@@
- Note on the assumptions: Prior work on CFG has shown that the linearity assumption holds in practice for low to moderate $w$ values (zheng2023characteristic). Furthermore, our use of an SEDD to sample $w$ during training (see Sec. 3 Incorporating Controllability) helps maintain accurate estimation of $\epsilon$ in the unconditional setting. Based on this, we conjecture that both assumptions approximately hold true in practice.
@@
Since the training loss is a squared error between two linear functions of $w$, minimizing this loss for all values of $w$ is equivalent to matching the corresponding coefficients of both functions. This means that to minimize the loss over the full support of $w$, the model must satisfy:
@@
$\epsilon_{\theta}(x_{t}\mid c,w,\mathcal{D}^{\mathcal{T}}$ $)=$ (21)
$\epsilon+(w-1)\cdot\left(\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})-\epsilon_{\theta_{0}}(x_{t})\right),$
which matches the DoG direction used at test time. Therefore, DogFit learns to internalize the DoG guidance signal during training, removing the need for external guidance at sampling. $\blacksquare$
@@
This enables DogFit to retain MG’s class-conditional guidance while avoiding out-of-distribution generation by pulling samples the target domain manifold’s core.
@@
According to Eq. 12 and Eq. 13, Eq. 24 can be rewritten:
$\epsilon$ $^{\prime}_{\text{{DogFit}}}-\epsilon^{\prime}_{\text{MG}}=$ (25)
$-\sigma(w-1)\cdot\text{sg}\left(\nabla_{x_{t}}\log p_{\theta}(x_{t}|\mathcal{D^{T}})-\nabla_{x_{t}}\log p_{\theta}(x_{t})\right).$
@@
Thus, given that MG uses $\epsilon_{\theta}(x_{t}|\mathcal{D^{T}})$ as an approximation of the term $\sigma_{t}\nabla_{x_{t}}\log p(x_{t}|\mathcal{D^{T}})$, we can express DogFit as applying the MG’s direction, but with a correction term that further pushes the sample towards a target domain classifier’s gradient direction. $\blacksquare$
@@ Appendix B Additional Experiments
We further examine the sensitivity of our method to $w$ in the absence of controllability. We train a specific model for each $w$ and compare the FID trends of DogFit and DoG across varying values (Fig. 10). While both methods achieve optimal FID in a similar range, DogFit exhibits a sharper degradation for larger $w$, suggesting higher sensitivity to strong guidance values. We hypothesize that this is due to training-time overexposure to extreme guidance, which can lead the model to learn from unnatural signals that do not follow the real target distribution, thereby reducing generation quality. This motivates our choice of sampling $w$ from an exponentially decaying distribution, ensuring the model is only sparsely exposed to high guidance values during training.
@@
Fig. 8 provides an extended version of the controllable guidance analysis for the Stanford-Cars dataset. This complements the main findings and illustrates the same trends on a new domain.
Metric Method Unlabelled Labelled
FFHQ ArtBench Caltech CUB-Birds Food Stanford-Cars Avg.
Precision (Inception) Fine-tuning 0.70 0.67 0.60 0.58 0.79 0.51 0.63
+ CFG (ho2022classifier) - 0.68 0.76 0.72 0.82 0.63 0.72
+ DoG (zhong2025domain) 0.83 0.77 0.81 0.82 0.88 0.77 0.81
MG (tang2025diffusion) - 0.68 0.76 0.71 0.81 0.62 0.72
[…]
Precision (DINOV2) Fine-tuning 0.74 0.84 0.48 0.16 0.41 0.30 0.44
+ CFG (ho2022classifier) - 0.83 0.67 0.30 0.48 0.47 0.52
+ DoG (zhong2025domain) 0.82 0.91 0.73 0.39 0.59 0.61 0.65
MG (tang2025diffusion) - 0.84 0.63 0.28 0.51 0.44 0.54
[…]
Recall (Inception) Fine-tuning 0.67 0.57 0.75 0.81 0.52 0.58 0.65
+ CFG (ho2022classifier) - 0.56 0.66 0.76 0.53 0.59 0.62
+ DoG (zhong2025domain) 0.61 0.50 0.65 0.68 0.48 0.52 0.57
MG (tang2025diffusion) - 0.56 0.67 0.75 0.55 0.59 0.624
[…]
Recall (DINOV2) Fine-tuning 0.46 0.21 0.66 0.95 0.67 0.82 0.66
+ CFG (ho2022classifier) - 0.22 0.59 0.87 0.61 0.84 0.63
+ DoG (zhong2025domain) 0.44 0.21 0.57 0.83 0.64 0.83 0.62
MG (tang2025diffusion) - 0.22 0.60 0.89 0.60 0.84 0.63
[…]
@@
Guidance Scheduling on $\text{FD}_{\text{DINOV2}}$
We observe that delaying guidance using the late-start strategy improves both $\text{FD}_{\text{DINOV2}}$ and precision, consistent with our FID-based findings. This supports the idea that postponing guidance allows for stronger and more accurate directional signals once the model has sufficiently stabilized (Fig. 9a). However, unlike FID, applying the cut-off strategy (limiting guidance to later denoising steps) shows little to no improvement in FDDINOV2 across most domains (Fig. 9 (b)). We believe this is due to a fundamental difference in what the two metrics capture: FID benefits from preserving the global structure inherited from the source model, which is often formed early in the denoising process. In contrast, FDDINOV2 focuses more on semantic alignment with the target domain, which may require guidance to be present throughout the generation, not just in the later steps. This distinction further highlights the importance of using multiple metrics when evaluating generative models.
Qualitative Comparison on SiT-XL/2
Fig. 11 provides a qualitative comparison on the SiT-XL/2 backbone. DogFit generates sharp, domain-relevant details on par with DoG, while avoiding out-of-distribution behavior occasionally seen with CFG and MG (e.g., distorted faces and birds).
Precision and Recall Analysis
Tab.4 presents a comparison of Precision and Recall (kynkaanniemi2019improved) between DogFit and the baselines on the DiT-XL/2 backbone. Precision reflects the fidelity or quality of generated samples, while Recall measures sample diversity. As expected, all guidance methods consistently improve Precision compared to fine-tuning while reducing Recall, aligning with the general behavior of guided diffusion models. This effect arises because guidance steers sampling toward high-likelihood regions of the conditional distribution, which inherently narrows the output distribution and suppresses diversity. Our method, both with and without guidance control, matches the performance of DoG in terms of both Precision and Recall in most settings. Notably, DogFit achieves the highest or second-highest Precision in both Inception and DINOV2 metrics across all datasets). In terms of Recall, DogFit and DoG slightly under-perform CFG and MG, which we believe is due to the additional domain alignment term further reducing the diversity to prevent out-of-domain target generations (Eq. 22). While recent works explore diversity-preserving techniques and scheduling mechanisms for guidance during sampling (kynkaanniemi2024applying; sadat2023cads), analogous techniques for training-time guidance are yet to be developed. We leave the exploration of such methods for DogFit as an exciting direction for future work.
@@
Tab.6 provides a comparison between DogFit and CFG when fine-tuning DiT-XL/2 pretrained at $512{\times}512$ on Food101 (bossard2014food) at $512{\times}512$. Results indicate that DogFit is fully scalable to larger resolutions, broadening its applicability and demonstrating invariance to image scale. For this experiment, we use a batch size of 12 and use the same hyperparameters as in our main experiments.
Robustness to Stronger Domain Shifts
While our target datasets span a range of different distribution gaps, we additionally evaluate transfer on the EuroSat dataset (helber2019eurosat), consisting of 27,000 satelite images with 10 classes, which shares very few features with the source domain. This setting probes whether DogFit’s source prior still provides a better signal than that of the fine-tuned target under extreme shift. We note that as the target domain diverges substantially from the source, the advantages of transfer learning from the source generally diminish relative to training on the target data from scratch. Nevertheless, our results in Tab.7 show that DogFit still outperforms Fine-tuning and CFG in this case, indicating that leveraging the source prior can still improve generation quality even when transfer benefits are attenuated.
@@
Tab.7 shows that sampling the guidance weight $\omega$ from a SEDD (Fig. 7) consistently outperforms uniform schedules across all test-time values of $\omega$. In particular, SEDD yields lower FID while uniform sampling over broader ranges destabilizes learning and collapses diversity, with FID inflating as $\omega$ increases. This empirical trend aligns with our linearity assumption in Prop, 1 which hints at the equivalence of DogFit and DoG (zhong2025domain) under small $\omega$ values (zheng2023characteristic). By concentrating the probability mass near $\omega\!\approx\!1$, SEDD keeps optimization within the proposition’s validity region, preserving the expected guidance behavior and allowing the use of larger test-time guidance values during generation.
DogFit in Text-to-Image Transfer Learning
To further show the broad applicability of our method, we test DogFit in two different text-to-image transfer settings, Dreambooth subject driven transfer  (ruiz2023dreambooth), and LoRA style transfer  (hu2022lora) to Pokémon. Note that these settings typically require large guidance values of up to $7.5$ for the model to produce visually pleasing images. However, DogFit is more theoretically grounded when exposed to small-guidance values during training, where it locally matches DoG (zhong2025domain) (see Prop. 1). While controllable guidance (conditioning on $w$) is a natural extension allowing for large test-time $\omega$ values, we focus here on fixed guidance during training and leave that extension for future work. Despite this conservative setting, DogFit significantly improves unguided fine-tuning with small guidance values ($\omega=[1.65,1.75]$) and approaches the performance of CFG with high guidance value ($\omega=7.5$).
@@
We apply Dreambooth fine-tuning on Stable Diffusion v1.5 (rombach2022high; podell2023sdxl) to novel subjects using a handful of instance images with DogFit. In this regime, the fine-tuning prompts are typically limited to templates such as “a [v] photo of {subject}, while at inference we expect the model to respond to diverse prompts from the source domain. This mismatch makes faithful distillation of text-conditioning challenging. Nevertheless, Fig. 12 and Tab. 8 show that DogFit learns controllable alignment with the prompt and significantly outperforms unguided fine-tuning while approaching CFG on DINO, CLIP-I, and CLIP-T, the standard metrics for subject-driven generation (ruiz2023dreambooth).
@@
We adapt Stable Diffusion v1.5 to the Pokémon style dataset [^1] https://huggingface.co/datasets/diffusers/pokemon-gpt4-captions with 833 image–caption pairs. We train a LoRA adapter (rank 4) for 2,000 iterations at a learning rate of $10^{-4}$. We evaluate with CLIP-T over the first 100 prompts (one image per prompt). As summarized in Tab.9, DogFit significantly improves prompt alignment over unguided fine-tuning while using a small guidance value ($\omega{=}1.75$) and a single sampling pass, and it closely approaches CFG which requires two passes and a large guidance value ($\omega{=}7.5$).